Use the properties of sets to prove that for all the sets A and B, A – (A ∩ B) = A – B

Grade:12th pass

1 Answers

Harshit Singh
3 years ago
Welcome to AskIITians

A – (A ∩ B) = A ∩ (A ∩ B)′ (since A – B = A ∩ B′)

= A ∩ (A′ ∪ B′) [by De Morgan’s law)

= (A∩A′) ∪ (A∩ B′) [by distributive law]

= φ ∪ (A ∩ B′)

= A ∩ B′ = A – B

Hence, proved that A – (A ∩ B) = A – B.

Thanks